The ratio of the measures of a pair of sides of a rectangle is 7:12. If the perimeter of the rectangle is 95, what is the length of each side?
Answer by stanbon(75887)

The ratio of the measures of a pair of sides of a rectangle is 7:12. If the perimeter of the rectangle is 95, what is the length of each side?
-------
Let width = 7x ; Then length = 12x
-----
P = 2(length + width)
----
95 = 2(12x + 7x)
19x = 95/2
x = 95/38
---
x = 2.5
----
width = 7x = 7*2.5 = 17.5
-----
length = 12x = 12*2.5 = 30
================================
Cheers,
Stan
============